OpenStudy (anonymous):

Tonya took out a loan to help pay for her house. She borrowed $60,000 for 15 years at a yearly simple interest rate of 5%. How much interest will she end up paying the bank? $450,000 $450 $45,000 $9,000

OpenStudy (anonymous):

@gigi2002 @iGreen.

OpenStudy (anonymous):

i think c

OpenStudy (anonymous):

it is c because you have to multiply all of the numbers. when you do that you end up with c.

OpenStudy (anonymous):

Maybe C is correct..I'm not sure. @ganeshie8 @dan815 @waterineyes I don't know how to do this problem, can any of you teach me?

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

it was c thanks

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

\[Simple \; \; Interest = \frac{Principle \times Rate \times Time}{100}\]

OpenStudy (anonymous):

August hosted two dinner parties for his friends. Twenty guests attended the first party, and twenty-nine guests attended the second party. What is the percentage increase of the number of guests from the first party to the second party? 45% 30% 55% 70% this one i think b

OpenStudy (anonymous):

Time should be in years..

OpenStudy (anonymous):

thx guys

OpenStudy (anonymous):

Are you getting @iGreen. ??

OpenStudy (anonymous):

Oh.. So it's (60,000 * 15 * 0.05)/100 ?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Do not convert Rate into fractions.. Rate = 5..

OpenStudy (anonymous):

Oh..it's just 5?

OpenStudy (anonymous):

Ah, I see

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

Getting 45,000??

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

\(\dagger\)

OpenStudy (anonymous):

Or we can just convert the rate to a decimal right?

OpenStudy (anonymous):

And then multiply..'cause I get 45,000 too 60,000 * 15 * 0.05

OpenStudy (anonymous):

yes, you can do so.. :)

OpenStudy (anonymous):

Oh, okay..thanks a bunch. I just medal'ed you and @CRAZYRVBGRIF

OpenStudy (anonymous):

@iGreen. ^_^ can u check my newest question

OpenStudy (anonymous):